### v5

@@ 4.5 Pre-trained Model Selection (RQ1)
Methodology. We shortlist the ST models for investigation in our work based on the NLP leaderboard, which reports the 38 most accurate pre-trained models [^8] https://www.sbert.net/docs/pretrained_models.html. These models have been extensively evaluated for their ability to generate sentence embeddings (i.e., capturing the semantics of the whole text) and their performance in semantic search (i.e., finding relevant answers to a given query). Both tasks closely align with our objectives. To identify trace links, we apply the pre-trained models in a zero-shot setting as follows. We let each model compute the similarity matrix equivalent to the output of step 5 in our approach (see Fig. 2). We then predict a trace link if the similarity value exceeds a predefined threshold. Since zero-shot does not require training, we simply run the pre-trained model on the entire HIPAA dataset.
@@
Evaluation Metrics. To better assess the performance irrespective of the selected threshold, we compute the Area Under the Curve (AUC) for the receiver operating characteristic (ROC) across different threshold values, ranging from $0.1$ to $0.9$. The ROC curve captures the trade-off between the true positive rate (TPR) and the false positive rate (FPR). TPR is the proportion of positives correctly identified as such (i.e., the percentage of trace links correctly identified for a given threshold). FPR is the proportion of negatives incorrectly identified as positives (i.e., the percentage of trace links wrongly identified as not trace links). The AUC of the ROC curve (computed as micro-average over all the provisions to avoid the dominance of some provisions) provides a single aggregate performance measure across all possible thresholds and, hence, is a suitable evaluation metric to compare the ST models. We posit that the model with the highest AUC value demonstrates the best overall performance in identifying trace links in a zero-shot setting, as a higher AUC value indicates a better balance between correctly identifying true trace links (high TPR) and minimizing the identification of false links (low FPR).
@@
The best-performing model on HIPAA is ST29 ($K^{\dagger}=1$), with an AUC value of 0.859. The next best performing model is ST21 with an AUC value of 0.850. The difference between these two AUC values is only marginal. A possible explanation is that ST29 uses ST21 as its base model. ST29 has been, however, trained on more (multi-lingual) data.
Additionally, we observe a discrepancy in the performance of the models on the HIPAA dataset compared to that reported by the NLP community. The best NLP model, ST1, does not perform well on HIPAA, ranked 16. This observation indicates that well-performing models in NLP are not necessarily as effective for RE-specific problems. We acknowledge that zero-shot performance does not always predict fine-tuning effectiveness, particularly across different model architectures. To address this concern, we fine-tuned the top five models based on their zero-shot AUC scores to evaluate their performance more comprehensively. However, our experiments show that ST29 still consistently outperforms the other models after fine-tuning. While resource constraints prevented us from fine-tuning all possible models, this targeted evaluation provides evidence that ST29 is the most effective choice among the top-performing candidates.
@@ 4.6 Accuracy on Benchmark Dataset (RQ2)
Methodology. We compare the four variants of $\mathcal{K}$ashif (explained in Section 3) against LSI, LDA, B from the literature cleland:2010 ; Guo:17 , GloVe, TraceBERT, RoBERTa, and LLaMa. We answer RQ2 on the benchmark dataset, HIPAA. Since HIPAA contains 10 requirements documents, we apply the leave-one-out (LOO) evaluation method, where $\mathcal{K}$ashif and baselines are tested on a left-out document each time and trained (or fine-tuned) on the remaining documents to emulate realistic settings. However, to ensure a reasonable balance between the training and test sets, we exclude one document (CCHIT, labeled H10 in Table 1) from the LOO process, as it contains 1,064 requirements, thereby including more than half the dataset.
Evaluation Metrics. We evaluate the four variations of $\mathcal{K}$ashif and the baselines using precision (P), measuring how many trace links identified by the approach are correct; recall (R), measuring how many trace links in our ground truth are correctly identified by the approach; and F2 score, the harmonic mean of precision and recall, with a greater emphasis on recall. We report the mean and standard deviation across the nine documents. To ease comparisons of the models’ rankings, we also include Mean Average Precision (MAP) over requirements that have at least one trace link to a provision, which measures how well a model ranks relevant items higher than irrelevant ones. We also report the AUC value for the overall performance of the classifiers. Moreover, we report the AUC value for each classifier to better evaluate their performance in distinguishing positive and negative pairs across thresholds. Note that the MAP score is computed only for requirements containing at least one true positive, since requirements without any trace links provide no informative contribution to the MAP. In contrast, AUC is applied to all requirements as it allows for a more complete assessment of each classifier’s capability to distinguish relevant from non-relevant pairs, regardless of the presence or absence of TPs for a particular requirement.
@@
Results. $\mathcal{K}$ashif $_{\text{tuned}}$ and the baseline models (LDA, LSI, B, and GloVe) require a decision threshold, which we tune using the same strategy explained in Section 3. In contrast, RoBERTa relies on the softmax function inherent to its architecture for final class-label assignment, and tuning the threshold over the output logits for label 1 resulted in a lower F2 score. Similarly, the publicly available implementations of TraceBERT and LLaMA output both class probabilities and a final label, and using the label as the final output yielded a higher F2 score. Therefore, these models do not require separate threshold selection. To compute the MAP value for RoBERTa, we use the output logits (scores) corresponding to the [CLS] token for label 1; for TraceBERT and LLaMa, we use the probability outputs.
@@
As visible from Table 5, $\mathcal{K}$ashif $_{\text{constant}}$ has the highest F2 score compared to the baselines. $\mathcal{K}$ashif also has the highest MAP score (81.4%), indicating the best ranking performance: in top-k results, it has the most TPs among baselines. Moreover, $\mathcal{K}$ashif has the highest AUC (0.93), indicating greater ability to distinguish between positive and negative pairs. $\mathcal{K}$ashif $_{\text{tuned}}$ outperforms all variants of $\mathcal{K}$ashif in terms of precision, achieving an average of 68.1%. Based on our observations of the tuned thresholds, we found that most threshold values are greater than 0.5, which explains the higher precision compared to $\mathcal{K}$ashif $_{\text{constant}}$. For the H1 to H9 projects, the corresponding threshold values are 0.91, 0.5, 0.79, 0.46, 0.82, 0.81, 0.78, 0.73, and 0.58, respectively.
The IR-based methods, LSI and LDA, exhibit notably different behaviors. LDA achieves an F2 score of only 14.4% and a MAP of 37.8%, indicating its failure to assign higher ranks to true-positive pairs. Furthermore, its AUC of 0.51 suggests that LDA performs no better than random and offers little discriminatory power. In contrast, LSI attains an F2 score of 10.7%, which is 4 pp lower than LDA. However, it outperforms LDA in both MAP (53.5%) and AUC (0.71). This suggests that LSI has stronger ranking capability. Additionally, Figure 5(b) shows that LSI’s maximum achievable F2 score is close to 20%, whereas LDA’s peak F2 score is around 14%. Therefore, the threshold-tuning process did not yield the best LSI performance.
B, the ML-based technique tailored for the LRT task, exhibits relatively low retrieval performance compared to the other classifiers. Nevertheless, it achieves the second-highest MAP score (78.8%), following $\mathcal{K}$ashif, indicating a strong ability to rank positive pairs higher than negative ones. We recall that B is a classifier that primarily uses a probabilistic approach based on word co-occurrence in requirement texts to predict whether each requirement should be traced to a particular provision. This indicates that B has greater discriminative ability between positive and negative pairs than IR-based techniques, GloVe, TraceBERT, and LLaMa. Moreover, Figure 5(b) shows that B’s maximum achievable F2 score is around 38%, indicating that threshold tuning was unable to identify an effective decision boundary for model B.
@@
TraceBERT, however, performs poorly on the LRT task, with results comparable to LDA. It achieves an F2 score of 8.7% and an AUC of 0.67, results that are only marginally better than random. However, its relatively high MAP of 62.7% indicates that it learned the relative similarity between pairs. We observed that TraceBERT produces nearly identical similarity values for most requirement-provision pairs, around 0.5. As a result, its threshold-based classification is substantially hindered. Moreover, in Figure 5(b), TraceBERT’s F2 score starts to change at a threshold value around 0.5. Recall that TraceBERT was originally designed for issue–code traceability and incorporates an intermediate training stage on a large external dataset. Therefore, its inability to achieve high performance in our setting is not unexpected.
RoBERTa achieves an F2 score of 41%, a MAP of 55.9%, and an AUC of 0.88, the second-highest AUC among all classifiers. Its high AUC value indicates that RoBERTa effectively separates positive and negative pairs. However, its ranking quality shows greater variability than $\mathcal{K}$ashif and B, as reflected in its MAP, which is fourth-highest overall. RoBERTa exhibits an almost constant F2 score across different threshold values, as shown in Figure 5(b). The maximum attainable F2 score for RoBERTa is around 41–43%, suggesting that the softmax-based output layer is already calibrated to produce its best possible performance within this range. Among the baselines, RoBERTa achieves the highest F2 score. RoBERTa also has the highest recall compared to all models and $\mathcal{K}$ashif variations, except for $\mathcal{K}$ashif Δ. However, compared to $\mathcal{K}$ashif constant, it has a lower recall, which shows weaker effectiveness in distinguishing relevant links by retrieving many FPs (650 FP links more than $\mathcal{K}$ashif constant).
Regarding the use of LLaMa as a classifier, we observe that LLaMa-w/o performs poorly with the lowest F2 score. After fine-tuning, however, the model’s performance improves significantly, demonstrating that it has learned task-specific cues. LLaMa-w/o has the AUC value of 0.50, which shows a random behavior and lacks discrimination on the LRT task. However, LLaMa-tuned achieves an F2 score of 22.5%, an AUC of 0.80, and a MAP of 41.4%, indicating that while the model can distinguish between positive and negative pairs (as reflected in its relatively high AUC), its overall ranking and threshold-based performance remain limited. The low MAP value suggests that LLaMa-tuned struggles to consistently assign high similarity scores to all true-positive pairs, capturing only a subset of links while missing many others. Moreover, in Figure 5(b), the highest attainable F2 score for LLaMa-tuned is above 25% and lower than 30%.
@@
Comparing the four variants of $\mathcal{K}$ashif, our results show that $\mathcal{K}$ashif $_{\text{constant}}$ is the best performing variant in terms of F2, achieving an average score of 62.7%. This score provides a gain of 4.2 pp over $\mathcal{K}$ashif $_{\text{tuned}}$, 13.7 pp over $\mathcal{K}$ashif dynamic, and 34.4 pp over $\mathcal{K}$ashif Δ. In terms of recall, however, $\mathcal{K}$ashif Δ achieves the best value of 80%, 12.7% more than $\mathcal{K}$ashif $_{\text{constant}}$. This can be explained by the threshold adjustment method for $\mathcal{K}$ashif Δ. Recall from Section 3 that to determine the threshold above which a trace link is predicted, we look at the largest gap in similarity values between the requirement and the provisions. Once determined, $\mathcal{K}$ashif Δ will always predict at least one trace link for each requirement corresponding to the provision with the highest similarity value that exceeds this gap. Such a method and recall value can indeed be useful when building recommendation systems. However, they come at the cost of introducing more FPs (as evidenced by the low precision), requiring the human analyst to filter them out. Consequently, we select $\mathcal{K}$ashif $_{\text{constant}}$ as the best performing model for LRT.
@@
    Computing high similarity scores when there are no trace links. A majority of FPs (96/113 = 84.9%) are due to falsely predicting a trace link for those requirements that have no trace links in our ground truth. This case suggests that a binary classifier could help in reducing FPs by predicting whether a requirement should have a trace link or not in the first place. We have conducted several experiments around this hypothesis. While we observed fewer FPs with a binary classifier, the overall improvement was not significant; hence, we do not report it in this paper.
@@
The answer to RQ2 is that $\mathcal{K}$ashif yields the best accuracy on HIPAA when we apply a constant threshold value of 0.5. Specifically, $\mathcal{K}$ashif achieves an F2 score of $\approx$63%. Compared to the best existing baseline from the literature, $\mathcal{K}$ashif has a gain of about 34 pp in F2 score. Moreover, $\mathcal{K}$ashif achieves the highest MAP score of 0.81 and AUC value of 0.93 compared to the baseline classifiers.
@@ 4.7 Effectiveness of Classification (RQ3)
Methodology. In real-life scenarios, dealing with LRT involves navigating through many provisions, usually significantly more than 10, as in the simple HIPAA case. This inherent complexity is notable with the 26 provisions pertinent to software in the GDPR. Using the test documents described in Section 4.2, we evaluate and compare two models, namely ST29—the best pre-trained ST model selected in RQ1 and $\mathcal{K}$ashif $_{\text{constant}}$—the best $\mathcal{K}$ashif variant fine-tuned on HIPAA identified in RQ2. Note that we opted not to fine-tune $\mathcal{K}$ashif again on the new documents for three reasons. First, the documents are small and thus inadequate for meaningful training (or fine-tuning). Second, we aim to challenge existing solutions with a more realistic scenario and apply them to new, unseen documents. Finally, $\mathcal{K}$ashif is a similarity-based solution which has been exposed to both the LRT task as well as the regulatory domain (terminology) in the first fine-tuning on HIPAA. Therefore, another fine-tuning is less likely to have any additional value.
@@
We also want to evaluate, on average, how well the model performs for a single requirement. Thus, we report two additional metrics: (iv) Macro Recall (MR), presented in Equation 1, to assess how well the predicted provisions are actually traced to the requirement; (v) Cost, presented in Equation 2, to capture the effort required for an analyst to check the predicted results and filter out the FPs.
@@
At the trace link level, we report the total number of actual trace links, true positives (TP) and false positives (FP), recall (R), precision (P), and F2 score.
@@
Results. Table 7 shows the results for each approach across the test documents, both at the trace link level and at the requirement level. The table shows the number of requirements in each test document [^9] Note that we leave out five requirements from each document to enable fair comparison with the Rice_LRT approach presented in RQ4, which relies on few-shot learning., the number of predicted trace links (T), TPs, FPs, R, P, and F2. It also includes the number of requirements with exact matches, the number with partial matches, and the success rate. From the table, we observe that ST29 performs worse than $\mathcal{K}$ashif in terms of success rate. While ST29 achieves a higher number of exact matches, it produces fewer partial matches compared to $\mathcal{K}$ashif. Note that, for both techniques, exact matches often represent “no trace link”, i.e., not predicting any trace link for requirements that had no trace links according to our ground truth. This highlights ST29’s failure to produce a non-empty set of trace links, resulting in a lower success rate. Therefore, allowing more partial matches to achieve a higher success rate, even at the cost of fewer exact matches, ensures broader coverage of correct links. Indeed, the effort required for a requirements analyst to filter out FPs is relatively low compared to the consequences of a requirement not being traced to a complete set of predicted provisions. Therefore, achieving a higher number of partial matches, despite FPs, can be more beneficial than achieving a higher number of exact matches for supporting requirements-level analysis. Our results indicate that the ST pre-trained model (ST29) failed to automatically predict trace links in most cases, suggesting that it neither understood the LRT task nor the application domain. Moreover, ST29 has the Cost of 0.06% and a MR of 1.3% per requirement. This means that for a requirement, on average, the model retrieves 0.06% of the provisions while ensuring 1.3% of the actual trace links are presented in those retrieved links. $\mathcal{K}$ashif achieves a Cost of 4.7% while with the MR value of 12.4%. Fine-tuning the models improves performance on the test documents; however, the patterns, which are more complex, in the unseen test documents, are still not fully captured by $\mathcal{K}$ashif.
@@
To summarize, $\mathcal{K}$ashif consistently outperforms ST29 across all documents, with a notable difference in the number of partial matches. Results thus suggest that fine-tuning pre-trained models on a dedicated dataset is indeed necessary for the model to learn about the LRT task. However, while better than the pre-trained model, $\mathcal{K}$ashif shows the following limitations: 1) it does not provide a rationale behind selecting a trace link, except for the fact that semantic similarity exceeds a pre-defined threshold. This is expected to impede its practical use. 2) The average success rate achieved by $\mathcal{K}$ashif is about 44%, which is not particularly effective.
@@
The answer to RQ3 is that $\mathcal{K}$ashif outperforms ST29, demonstrating that fine-tuning helps the model learn about the LRT task. However, the performance of $\mathcal{K}$ashif still shows significant room for improvement in unseen domains.

### v6

@@ 4.5 Pre-trained Model Selection (RQ1)
Methodology. We shortlist the ST models for investigation in our work based on the NLP leaderboard, which reports the 38 most accurate pre-trained models [^8] https://www.sbert.net/docs/pretrained_models.html. These models have been extensively evaluated for their ability to generate sentence embeddings (i.e., capturing the semantics of the entire text) and for their performance in semantic search (i.e., retrieving relevant answers to a given query). Both tasks closely align with our objectives. To identify trace links, we apply the pre-trained models in a zero-shot setting as follows. We let each model compute the similarity matrix equivalent to the output of step 5 in our approach (see Fig. 2). We then predict a trace link if the similarity value exceeds a predefined threshold. Since zero-shot does not require training, we simply run the pre-trained model on the entire HIPAA dataset.
@@
Evaluation Metrics. To better assess the performance irrespective of the selected threshold, we compute the Area Under the Curve (AUC) for the receiver operating characteristic (ROC) across different threshold values, ranging from $0.1$ to $0.9$. The ROC curve captures the trade-off between the true positive rate (TPR) and the false positive rate (FPR). TPR is the proportion of positives correctly identified as such (i.e., the percentage of trace links correctly identified for a given threshold). FPR is the proportion of negatives incorrectly identified as positives (i.e., the percentage of trace links wrongly identified as not trace links). The AUC of the ROC curve (computed as a micro-average across all provisions to avoid the dominance of any single provision) provides a single aggregate performance measure across all possible thresholds and, hence, is a suitable evaluation metric for comparing ST models. We posit that the model with the highest AUC performs best in identifying trace links in a zero-shot setting, as a higher AUC indicates a better balance between correctly identifying true trace links (high TPR) and minimizing false positives (low FPR).
@@
The best-performing model on HIPAA is ST29 ($K^{\dagger}=1$), with an AUC value of 0.859. The next best-performing model is ST21, with an AUC of 0.850. The difference between these two AUC values is only marginal. A possible explanation is that ST29 uses ST21 as its base model. ST29 has been, however, trained on more (multi-lingual) data.
Additionally, we observe a discrepancy between the models’ performance on the HIPAA dataset and that reported by the NLP community. The best NLP model, ST1, does not perform well on HIPAA, ranking 16th. This observation indicates that high-performing NLP models are not necessarily effective for RE-specific problems. We acknowledge that zero-shot performance does not always predict the effectiveness of fine-tuning, particularly across different model architectures. To address this concern, we fine-tuned the top five models based on their zero-shot AUC scores to evaluate their performance more comprehensively. However, our experiments show that ST29 continues to outperform the other models after fine-tuning. While resource constraints prevented us from fine-tuning all possible models, this targeted evaluation provides evidence that ST29 is the most effective choice among the top-performing candidates.
@@ 4.6 Accuracy on Benchmark Dataset (RQ2)
Methodology. We compare the four variants of $\mathcal{K}$ashif (explained in Section 3) against LSI, LDA, B from the literature cleland:2010 ; Guo:17 , GloVe, TraceBERT, RoBERTa, and LLaMa. We answer RQ2 on the benchmark dataset, HIPAA. Since HIPAA comprises 10 requirements documents, we apply the leave-one-out (LOO) evaluation method, in which $\mathcal{K}$ashif and baselines are tested on a left-out document each time and trained (or fine-tuned) on the remaining documents to emulate realistic settings. However, to ensure a reasonable balance between the training and test sets, we exclude one document (CCHIT, labeled H10 in Table 1) from the LOO process because it contains 1,064 requirements, thereby comprising more than half of the dataset.
Evaluation Metrics. We evaluate the four variations of $\mathcal{K}$ashif and the baselines using precision (P), measuring how many trace links identified by the approach are correct; recall (R), measuring how many trace links in our ground truth are correctly identified by the approach; and F2 score, the harmonic mean of precision and recall, with a greater emphasis on recall. We report the mean and standard deviation across the nine documents. To facilitate comparisons of the models’ rankings, we also report Mean Average Precision (MAP) over requirements with at least one trace link to a provision, which measures how well a model ranks relevant items higher than irrelevant ones. We also report the AUC value for the overall performance of the classifiers.
Moreover, we report the AUC value for each classifier to better evaluate their performance in distinguishing positive and negative pairs across thresholds. Note that the MAP score is computed only for requirements containing at least one true positive, since requirements without any trace links provide no informative contribution to the MAP. In contrast, AUC is applied to all requirements because it enables a more comprehensive assessment of each classifier’s ability to distinguish relevant from non-relevant pairs, regardless of the presence or absence of TPs for a particular requirement.
@@
Results. $\mathcal{K}$ashif $_{\text{tuned}}$ and the baseline models (LDA, LSI, B, and GloVe) require a decision threshold, which we tune using the same strategy explained in Section 3. In contrast, RoBERTa relies on the softmax function inherent to its architecture for final class-label assignment, and tuning the threshold on the output logits for label 1 yielded a lower F2 score. Similarly, the publicly available implementations of TraceBERT and LLaMA output both class probabilities and a final label, and using the label as the final output yielded a higher F2 score. Therefore, these models do not require separate threshold selection. To compute the MAP value for RoBERTa, we use the output logits (scores) corresponding to the [CLS] token for label 1; for TraceBERT and LLaMa, we use the probability outputs.
@@
As visible from Table 5, $\mathcal{K}$ashif $_{\text{constant}}$ has the highest F2 score compared to the baselines. $\mathcal{K}$ashif also has the highest MAP score (81.4%), indicating the best ranking performance: in top-k results, it has the most TPs among baselines. Moreover, $\mathcal{K}$ashif has the highest AUC (0.93), indicating greater ability to distinguish between positive and negative pairs. $\mathcal{K}$ashif $_{\text{tuned}}$ outperforms all variants of $\mathcal{K}$ashif in terms of precision, achieving an average of 68.1%. Based on our observations of the tuned thresholds, we found that most threshold values exceed 0.5, which explains the higher precision relative to $\mathcal{K}$ashif $_{\text{constant}}$. For the H1 to H9 projects, the corresponding threshold values are 0.91, 0.5, 0.79, 0.46, 0.82, 0.81, 0.78, 0.73, and 0.58, respectively.
The IR-based methods, LSI and LDA, exhibit notably different behaviors. LDA achieves an F2 score of only 14.4% and a MAP of 37.8%, indicating its failure to assign higher ranks to true-positive pairs. Furthermore, its AUC of 0.51 indicates that LDA performs no better than chance and provides little discriminatory power. In contrast, LSI achieves an F2 score of 10.7%, which is 4 pp lower than that of LDA. However, it outperforms LDA in both MAP (53.5%) and AUC (0.71). This suggests that LSI has stronger ranking capability. Additionally, Figure 5(b) shows that LSI’s maximum achievable F2 score is close to 20%, whereas LDA’s peak F2 score is around 14%. Therefore, the threshold-tuning process did not yield the best LSI performance.
B, an ML-based technique tailored to the LRT task, exhibits relatively low retrieval performance compared with the other classifiers. Nevertheless, it achieves the second-highest MAP score (78.8%), following $\mathcal{K}$ashif, indicating a strong ability to rank positive pairs higher than negative ones. We recall that B is a classifier that primarily uses a probabilistic approach based on word co-occurrence in requirement texts to predict whether each requirement should be traced to a particular provision. This indicates that B has greater discriminative ability between positive and negative pairs than IR-based techniques, GloVe, TraceBERT, and LLaMa. Moreover, Figure 5(b) shows that B’s maximum achievable F2 score is around 38%, indicating that threshold tuning was unable to identify an effective decision boundary for model B.
@@
TraceBERT, however, performs poorly on the LRT task, yielding results comparable to those of LDA. It achieves an F2 score of 8.7% and an AUC of 0.67, results that are only marginally better than random. However, its relatively high MAP of 62.7% indicates that it learned the relative similarity between pairs. We observed that TraceBERT produces nearly identical similarity values for most requirement-provision pairs, around 0.5. As a result, its threshold-based classification is substantially hindered. Moreover, in Figure 5(b), TraceBERT’s F2 score starts to change at a threshold value around 0.5. Recall that TraceBERT was originally designed for issue–code traceability and incorporates an intermediate training stage on a large external dataset. Therefore, its inability to achieve high performance in our setting is not unexpected.
RoBERTa achieves an F2 score of 41%, a MAP of 55.9%, and an AUC of 0.88, the second-highest AUC among all classifiers. Its high AUC value indicates that RoBERTa effectively separates positive and negative pairs. However, its ranking quality shows greater variability than that of $\mathcal{K}$ashif and B, as reflected in its MAP, which is the fourth-highest overall. RoBERTa exhibits an almost constant F2 score across different threshold values, as shown in Figure 5(b). The maximum attainable F2 score for RoBERTa is around 41–43%, suggesting that the softmax-based output layer is already calibrated to produce its best possible performance within this range. Among the baselines, RoBERTa achieves the highest F2 score. RoBERTa also achieves the highest recall among all models and $\mathcal{K}$ashif variations, except for $\mathcal{K}$ashif Δ. However, compared to $\mathcal{K}$ashif constant, it has a lower recall, which shows weaker effectiveness in distinguishing relevant links by retrieving many FPs (650 FP links more than $\mathcal{K}$ashif constant).
Regarding the use of LLaMa as a classifier, we observe that LLaMa-w/o performs poorly with the lowest F2 score. After fine-tuning, however, the model’s performance improves significantly, demonstrating that it has learned task-specific cues. LLaMa-w/o has the AUC value of 0.50, which shows a random behavior and lacks discrimination on the LRT task. However, LLaMa-tuned achieves an F2 score of 22.5%, an AUC of 0.80, and a MAP of 41.4%, indicating that, although the model can distinguish positive from negative pairs (as reflected in its relatively high AUC), its overall ranking and threshold-based performance remain limited. The low MAP value suggests that LLaMa-tuned struggles to consistently assign high similarity scores to all true-positive pairs, capturing only a subset of links while missing many others. Moreover, in Figure 5(b), the highest attainable F2 score for LLaMa-tuned is above 25% and lower than 30%.
@@
Comparing the four variants of $\mathcal{K}$ashif, our results show that $\mathcal{K}$ashif $_{\text{constant}}$ is the best performing variant in terms of F2, achieving an average score of 62.7%. This score provides a gain of 4.2 pp over $\mathcal{K}$ashif $_{\text{tuned}}$, 13.7 pp over $\mathcal{K}$ashif dynamic, and 34.4 pp over $\mathcal{K}$ashif Δ. In terms of recall, however, $\mathcal{K}$ashif Δ achieves the best value of 80%, 12.7% more than $\mathcal{K}$ashif $_{\text{constant}}$. This can be explained by the threshold adjustment method for $\mathcal{K}$ashif Δ. Recall from Section 3 that to determine the threshold above which a trace link is predicted, we look at the largest gap in similarity values between the requirement and the provisions. Once determined, $\mathcal{K}$ashif Δ will always predict at least one trace link for each requirement corresponding to the provision with the highest similarity value that exceeds this gap. Such a method and recall value can indeed be useful when building recommendation systems. However, they come at the cost of increasing FPs (as evidenced by low precision), requiring the human analyst to filter them out. Consequently, we select $\mathcal{K}$ashif $_{\text{constant}}$ as the best performing model for LRT.
@@
    Computing high similarity scores when there are no trace links. A majority of FPs (96/113 = 84.9%) are due to falsely predicting a trace link for those requirements that have no trace links in our ground truth. This case suggests that a binary classifier could help in reducing FPs by predicting whether a requirement should have a trace link or not in the first place. We have conducted several experiments around this hypothesis. Although we observed fewer FPs with a binary classifier, the overall improvement was not statistically significant; therefore, we do not report it in this paper.
@@
The answer to RQ2 is that $\mathcal{K}$ashif yields the best accuracy on HIPAA when we apply a constant threshold value of 0.5. Specifically, $\mathcal{K}$ashif achieves an F2 score of $\approx$63%. Compared with the best existing baseline in the literature, $\mathcal{K}$ashif achieves a gain of approximately 34 points in F2 score. Moreover, $\mathcal{K}$ashif achieves the highest MAP score of 0.81 and AUC value of 0.93 compared to the baseline classifiers.
@@ 4.7 Effectiveness of Classification (RQ3)
Methodology. In real-life scenarios, dealing with LRT involves navigating through many provisions, usually significantly more than 10, as in the simple HIPAA case. This inherent complexity is notable with the 26 provisions pertaining to software in the GDPR. Using the test documents described in Section 4.2, we evaluate and compare two models, namely ST29—the best pre-trained ST model selected in RQ1 and $\mathcal{K}$ashif $_{\text{constant}}$—the best $\mathcal{K}$ashif variant fine-tuned on HIPAA identified in RQ2. Note that we opted not to fine-tune $\mathcal{K}$ashif again on the new documents for three reasons. First, the documents are too small to support meaningful training (or fine-tuning). Second, we aim to evaluate existing solutions in a more realistic scenario and apply them to unseen documents. Finally, $\mathcal{K}$ashif is a similarity-based solution that has been exposed to both the LRT task and the regulatory domain (terminology) during the first fine-tuning on HIPAA. Therefore, further fine-tuning is unlikely to add value.
@@
We also aim to evaluate, on average, the model’s performance for a single requirement. Thus, we report two additional metrics: (iv) Macro Recall (MR), presented in Equation 1, to assess how well the predicted provisions are actually traced to the requirement; (v) Cost, presented in Equation 2, to capture the effort required for an analyst to check the predicted results and filter out the FPs.
@@
At the trace-link level, we report the total number of trace links, true positives (TPs), false positives (FPs), recall (R), precision (P), and F2 score.
@@
Results. Table 7 shows the results for each approach across the test documents, both at the trace link level and at the requirement level. The table shows the number of requirements in each test document [^9] Note that we leave out five requirements from each document to enable fair comparison with the Rice_LRT approach presented in RQ4, which relies on few-shot learning., the number of predicted trace links (T), TPs, FPs, R, P, and F2. It also includes the number of requirements with exact matches, the number with partial matches, and the success rate. From the table, we observe that ST29 performs worse than $\mathcal{K}$ashif in terms of success rate. While ST29 achieves a higher number of exact matches, it produces fewer partial matches compared to $\mathcal{K}$ashif. Note that, for both techniques, exact matches often represent “no trace link”, i.e., not predicting any trace link for requirements that had no trace links according to our ground truth. This highlights ST29’s failure to produce a non-empty set of trace links, resulting in a lower success rate. Therefore, allowing more partial matches to achieve a higher success rate, even at the cost of fewer exact matches, ensures broader coverage of correct links. Indeed, the effort required for a requirements analyst to filter out FPs is relatively low compared to the consequences of a requirement not being traced to a complete set of predicted provisions. Therefore, achieving a higher number of partial matches, despite FPs, can be more beneficial than achieving a higher number of exact matches for supporting requirements-level analysis. Our results indicate that the ST pre-trained model (ST29) failed to automatically predict trace links in most cases, suggesting that it neither understood the LRT task nor the application domain. Moreover, ST29 has the Cost of 0.06% and a MR of 1.3% per requirement. This means that, on average, for a given requirement, the model retrieves 0.06% of the provisions while ensuring that 1.3% of the actual trace links are included among those retrieved. $\mathcal{K}$ashif achieves a Cost of 4.7% while with the MR value of 12.4%. Fine-tuning the models improves performance on the test documents; however, the more complex patterns in the unseen test documents are still not fully captured by $\mathcal{K}$ashif.
@@
To summarize, $\mathcal{K}$ashif consistently outperforms ST29 across all documents, with a notable difference in the number of partial matches. Results thus suggest that fine-tuning pre-trained models on a dedicated dataset is necessary for the models to learn the LRT task. However, while better than the pre-trained model, $\mathcal{K}$ashif shows the following limitations: 1) it does not provide a rationale behind selecting a trace link, except for the fact that semantic similarity exceeds a pre-defined threshold. This is expected to impede its practical use. 2) The average success rate achieved by $\mathcal{K}$ashif is about 44%, which is not particularly effective.
@@
The answer to RQ3 is that $\mathcal{K}$ashif outperforms ST29, demonstrating that fine-tuning helps the model learn about the LRT task. However, kashif’s performance still shows significant room for improvement in unseen domains.